OpenStudy (anonymous):

Two angles are complementary. The measure of the larger angle is five times the measure of the smaller angle. What is the measure of the larger angle?

OpenStudy (anonymous):

let x be the larger angle and y the smaller angle. so x=5y since larger angle is five times the measure of the smaller anlge when two angles are said to be complementary then it means that their sums must be equal to 90 degrees so x + y= 90. replace x=5y in x + y=90 hence 5y + y = 90 6y =90 y=15 and finally x=5y=5.(15)=75

OpenStudy (anonymous):

x=75 degrees is the measure of the larger angle